# Wenfan Ou and Sheng Bi

### v3

@@
Fueled by both advances in robust optimization theory and applications of reinforcement learning, robust Markov Decision Processes (RMDPs) have gained increasing attention, due to their powerful capability for sequential decision-making under uncertainty. This review provides an in-depth overview of the evolution and advances in RMDPs formulations, particularly in ambiguity modeling, and classifies these methods for representing uncertainty into three principal approaches: parametric, moment-based, and discrepancy-based, elaborating the trade-offs among the alternative representations. Meanwhile, the review delves into the rectangular assumptions, which guarantee the tractability of RMDPs yet are noted for their conservatism. The review summarizes three popular rectangular conditions and develops a new proof to attest to the NP-hardness of non-rectangular RMDPs. Out of the traditional RMDPs scope, recent efforts without conventional rectangular assumptions and new fashions within the RMDPs community are also reviewed. These studies foster the development of more flexible and practical modeling frameworks and enhance the adaptability and performance of RMDPs.
@@
robust Markov Decision Processes, ambiguity sets modeling, reinforcement learning, rectangularity, new fashions
@@ 1 Introduction
Markov decision processes (MDPs) originate from the seminal works of BELLMAN (1957) and Howard (1960) to model sequential decision-making problems under uncertainty. As powerful mathematical tools, MDPs are widely applied in many areas, such as economics, transportation, inventory control, health care, and manufacture (Parkes & Singh 2003, McKenna et al. 2020, Song & Zipkin 1993, Katehakis & Puranam 2012, Fan et al. 2022, Che et al. 2023). Recently, reinforcement learning (RL), which establishes its theoretical foundations on MDPs (Sutton & Barto 2018), has attracted increasing attention and achieved notable success in practical applications. RL stands out for its potential intelligence on decision-making and active learning capabilities, especially when incorporating deep learning (Matsuo et al. 2022, Kosanoglu et al. 2022, Bushaj et al. 2023, Rolf et al. 2023, Zuccotto et al. 2024). Undeniably, the advances in RL pave the way for wider applications of MDPs. However, volatile and dynamic environments always cause poor performance of MDPs for sequential decision-making. Naive MDPs presuppose a complete grasp of the underlying dynamics, an assumption that often falls short in the real world. With deviated estimation or belief affected by uncertainty, decision-makers should anticipate disappointment on average, owing to the optimization-based selection process. This is the so-called the optimizer’s curse (Smith & Winkler 2006). Even worse, this skewed belief in sequential decision-making will persist throughout the entire planning period, persistently leading to worse outcomes compared to single-stage decision-making problems.
Motivated by advances in robust optimization theory and practical performance (Scarf et al. 1957, Ben-Tal & Nemirovski 2002, Bertsimas & Sim 2004, Goerigk & Schöbel 2016), robust Markov decision processes (RMDPs) have been investigated for decades. Unlike conventional MDPs, RMDPs are specifically designed to deal with situations where only partial information about the true underlying parameters or probability distributions is available, rather than complete knowledge.
The decision-maker can construct a collection of possible parameter values or distributions, called an uncertainty set or ambiguity set, based on the belief that the partial information is credible. Subsequently, the decision-maker can determine the optimal solution under the worst-case scenario over the uncertainty or ambiguity set.
The exploration of RMDPs can be traced back to the 1970s (Satia & Lave Jr 1973) when they were known as MDPs with imprecisely known parameters (MDPIPs) (White III & El-Deib 1986, White III & Eldeib 1994, Givan et al. 2000). Most MDPIPs assume the transition probabilities are bounded by linear programs, appealing to linear programming at each iteration. While defining the uncertainty on the transition matrix by a polytope is natural and convenient, the approach brings some drawbacks. Polytopic models not only pose computational challenges but also lead to excessively conservative policies due to the possibly terrible representations of statistical uncertainty. To address this issue, seminar papers by Iyengar (2005) and Nilim & El Ghaoui (2005) independently propose a similar formulation that leverages strong duality proofs and rectangularity assumption to extend value iteration for MDPs to robust value iteration. This formulation is also referred to as robust DP, and serves as a modern incarnation of RMDPs thereafter. Both two papers introduce two imperative concepts. Firstly, they introduce ambiguity sets, especially $\phi$-divergences, norm, and likelihood region ambiguity sets, which outperform polytopic models. Secondly, these works coincidentally emphasize that the rectangularity assumption is essential for the tractability of RMDPs, albeit without providing concrete proof.
To fill the theoretical gap, Bagnell et al. (2001) and Wiesemann et al. (2013) provide the NP-hard proofs for general (non-rectangular) RMDPs in different ways, respectively. Moreover, with the conservatism of robust paradigm proved widely (Thiele 2010, Delage & Mannor 2010) and innovations of distributionally robust optimization (DRO) (Popescu 2007, Delage & Ye 2010, Goh & Sim 2010), Xu & Mannor (2012) first propose the framework of distributionally robust MDPs (DRMDPs). Compared to traditional RMDPs, the new framework is more convenient to incorporate statistical information to mitigate conservatism. Xu & Mannor (2012) and Yu & Xu (2015) show that DRMDPs can be reduced to expected standard RMDPs under some assumptions, thereby retaining the tractability. These research findings have laid a solid theoretical foundation for RMDPs, and enable the thriving development of the RMDP community.
Beneficial from substantial results across diverse fields and interests in data-driven practices, the literature on RMDPs has been significantly enriched in recent years. Researchers have explored RMDP modeling from various aspects, including the exploration of diverse ambiguity sets, the development of effective solving and approximation algorithms, as well as the design of robust algorithms for RL. Furthermore, RMDPs have been combined with other techniques from various perspectives and applied in many areas. These techniques include online learning (Croonenborghs et al. 2007, Badrinath & Kalathil 2021, Cowan et al. 2018), safe learning (Hans et al. 2008, Polo & Rebollo 2011, Wachi & Sui 2020), dynamic risk measure (Bäuerle & Rieder 2014, Chow et al. 2015, Yu & Shen 2022, Rockafellar & Uryasev 2000), and regularization (Farahmand 2011, Zhang et al. 2018), etc. Despite the surge in the literature on RMDPs in recent years, a systematic review of RMDPs formulation remains scarce, with only a handful of tutorials or general reviews available (Mannor & Xu 2019, Keith & Ahner 2021, Badings et al. 2023). We believe that the absence of a specialized review hinders the development of RMDP theory and applications. To bridge this gap, we provide an up-to-date review of the RMDPs/DRMDPs formulations with uncertain transitions. This is a common class of RMDPs problems that have been widely considered in research and applications within the management science and operations research community. By surveying impressive results and studies in this domain, we aim to establish a solid foundation for future research and practical implementations in RMDPs.
@@
The remainder of the paper is organized as follows. In Section 2, we briefly introduce the preliminary of RMDPs. In particular, we summarize the definitions of rectangularity and provide a new proof for NP-hard of non-rectangular RMDPs. In Section 3, 4 and 5, we respectively introduce three popular types of RMDPs: parametric RMDPs, moment-based RMDPs, and discrepancy-based RMDPs. In these sections, we focus on generic formulations and key theoretical advancements. Although rectangularity maintains the tractability of RMDPs, it often doesn’t hold in practice. In Section 6, we investigate methodologies for modeling coupled uncertainty that violates rectangularity assumptions. Finally, we review some novel works in Section 7 that have sparked our interest beyond the traditional mini-max framework and conclude in Section 8.
@@ 2 Preliminaries
We denote by $[T]=\{1,2,...,T\}$ the set of positive running indices up to $T$, and use $\bigotimes$ to represent the Cartesian product. Let $\mathfrak{M}(\Xi)$ be a set of probability measures on measurable space $(\Xi,\mathcal{F})$, where $\Xi:=\bigotimes_{t\in[T]}\Xi_{t}$ is finite, and $\Xi_{t}$ is the sample space at stage $t$. Define $\left\{\emptyset,\Xi\right\}=\mathcal{F}_{0}\subseteq\mathcal{F}_{1}\subseteq\cdots\subseteq\mathcal{F}_{T}=\mathcal{F}$. Let $\mathcal{P}\subseteq\mathfrak{M}(\Xi)$ be the set of probability measures we are interested in, and $P\in\mathcal{P}$ be a probability measure on $(\Xi,\mathcal{F})$. Without loss of generality, we can appropriately derive the corresponding probabilities $\mathbf{p}$ or a (discrete) distribution $\mu$ with respect to the probability measure $P$. Let $\mathcal{P}_{t}(\xi)=\left\{P(\cdot|\mathcal{F}_{t})(\xi)\ |\ P\in\mathcal{P}\right\}$. $\Delta^{d}=\left\{\mathbf{p}\ |\ p_{1}+\cdots+p_{d}=1,p_{i}\geq 0\ \forall i\in[d]\right\}\in\mathbb{R}^{d}$ is the probability simplex. $\bm{1}(\cdot)$ is an indicator function. We use superscript 00 to denote true underlying value, e.g., $\mathbf{p}^{0}$, and $\hat{\cdot}$ denotes the estimated or nominal one, e.g., $\hat{\mathbf{p}}$.
[…]
@@
where $0<\gamma\leq 1$ is the discounted factor. Consequently, the (randomized) policy $\pi:\mathcal{S}\rightarrow\Delta^{|\mathcal{A}|}$ specifies a decision-making strategy in which the decision-maker chooses actions adaptively based on the current state, i.e., takes action $a\in\mathcal{A}$ with probability $\pi(s,a)$ at state $s\in\mathcal{S}$. If $\pi:\mathcal{S}\rightarrow\mathcal{A}$, we call it a deterministic policy.
Let $V_{t}(\cdot):\mathcal{S}\rightarrow\mathbb{R}$ denote the value-to-go function at stage $t$, solving (1) is equivalent to solving the following equations recursively:
$V_{t}(s_{t})=\max_{\pi}\mathbb{E}_{\mathbf{p}}^{\pi}\left[R(s_{t},a_{t},s_{t+1})+\gamma V_{t+1}(s_{t+1})\right]$ (2)
where $V_{T+1}$ is the terminal value which takes the value of 0 or $-R(s_{T+1})$ conventionally.
The value-to-go function (2) can be solved efficiently by dynamic programming techniques (finite horizons) and value/policy iteration methods (infinite horizons) (Puterman 2014), although there exists the so-called “curse of dimensionality” (Bellman 1966).
Note that with the growing research on MDPs, there have been diverse representations of MDPs proposed, such as assuming known initial state distributions or continuous state spaces. However, the essential elements of MDPs outlined above capture the core components and remain widely applicable across this review, unless otherwise specified. Future extensions can build on this general MDP formulation by incorporating additional details and specifications tailored to particular problems.
@@
The transition probability matrix $\mathbf{p}$ defined above is associated with the transition tuple $(s,a,s^{\prime})$, but most literature assumes exogenous randomness. For example, in inventory management problems, the demand is often not affected by inventory policies. Therefore, $\mathbf{p}$ may be independent of actions and can be represented by a distribution directly.
@@
As mentioned in Section 1, we focus on MDPs with uncertain transition probability. To ensure generality, we claim that the set $\mathcal{P}$ consists of probability measures on probability space. However, we sometimes abuse the notation $\mathcal{P}$ to represent the set of corresponding transition probabilities $\mathbf{p}$ or distribution $\mu$ to align with literature. Formally, we consider RMDPs and DRMDPs, defined as follows:
@@
A RMDP is defined as a tuple $<T,\mathcal{S},\mathcal{A},\mathcal{P},\gamma>$, where the transition probability matrix $\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}$ is unknown but belongs to a known collection $\mathcal{P}$ of parameters, called uncertainty set. Exactly, RMDPs can be expressed as
@@
A DRMDP can be represented by a tuple $<T,\mathcal{S},\mathcal{A},\mathcal{P},\gamma>$, where the transition probability matrix $\mathbf{p}$ obeys to a distribution $\mu$, denoted by $\mathbf{p}\sim\mu$. $\mu$ is unknown but belongs to a known distribution family $\mathcal{P}$ of probability distributions, called the ambiguity set. Exactly, DRMDPs can be expressed as
@@
While RMDPs view the transition probabilities as uncertain parameters, DRMDPs model them as random variables or use the distribution of exogenous randomness to represent dynamics. This difference in definitions is noteworthy, bringing distinct formulation styles and enabling distinct ambiguity set modeling. For example, moment-based ambiguity sets are more prevalent in DRMDPs compared to RMDPs.
Despite the differences between RMDPs and DRMDPs, they exhibit similar formulations within the scope of this article. The high-level modeling approaches exhibit substantial parallels, and both formulations can be converted to each other with appropriate modeling and practical undergrounds, especially in a discrepancy-based framework (Section 5). Thus, this article mainly investigates the literature based on modeling frameworks, rather than specifically differentiating RMDPs and DRMDPs. Unless otherwise noted, RMDPs and DRMDPs will both be referred to as RMDPs in the following for simplicity.

### v4

@@
Fueled by advances in both robust optimization theory and reinforcement learning (RL), robust Markov Decision Processes (RMDPs) have garnered increasing attention due to their powerful capability for sequential decision-making under uncertainty. In this paper, we provide a comprehensive overview of the theoretical foundations and recent developments in RMDPs, with a particular emphasis on ambiguity modeling. We examine the “rectangular assumption”, a key condition ensuring computational tractability in RMDPs but often resulting in overly conservative policies. Three widely used rectangular forms are summarized, and a novel proof is provided for the NP-hardness of non-rectangular RMDPs. We categorize RMDP formulation approaches into parametric, moment-based, and discrepancy-based models, analyzing the trade-offs associated with each representation. Beyond the traditional scope of RMDPs, we also explore recent efforts to relax rectangular assumptions and highlight emerging trends within the RMDP research community. These developments contribute to more practical and flexible modeling frameworks, complementing the classical RMDP results. Relaxing rectangular assumptions tailored to operations management is a promising area for future research, and there are also opportunities for further advances in developing fast algorithms and provably robust RL algorithms.
@@
robust Markov decision processes, ambiguity sets modeling, reinforcement learning, rectangularity
@@ 1 Introduction
[…]
In this survey, we focus on the problems where the transition probabilities are unknown. We assume rewards either entirely depend on the current state and action or are governed by the transition probabilities. Full knowledge of the transition probabilities is usually unavailable in the real world. Under a data-driven paradigm, the decision-maker can only estimate model parameters or distributions based on historical data, which introduces additional external uncertainty (distinguished from the internal uncertainty due to the stochastic nature of MDPs). With the deviated estimation or belief affected by external uncertainty, the decision-maker should anticipate disappointment on average, owing to the optimization-based selection process. This is the so-called the optimizer’s curse (Smith & Winkler 2006). Even worse, since sequential decision-making can amplify decision errors across stages, inaccurate estimations often result in significantly worse outcomes than in single-stage problems that involve only one decision point. We refer to the case study in Mannor et al. (2007) to demonstrate the influence of external uncertainty under the optimization process.
[…]
Regarding the computationally tractable recursive structure of MDP, another natural extension to multi-stage settings is robust MDP (RMDP). In recent years, driven by breakthroughs in theory and practical needs of data-driven paradigms, RO has seen substantial progress, and RMDP has also been emerging within operations research and computer science communities. Unlike conventional MDPs, RMDPs are tailored for circumstances where only partial information about the true underlying parameters or probability distributions is available, rather than complete knowledge. The decision-maker can construct a collection of possible parameter values (resp. distributions), called an uncertainty set (resp. ambiguity set), based on the belief that the partial information is credible. This allows the decision-maker to determine the optimal solution under the worst-case scenario over the uncertainty or ambiguity set.
[…]
To address this issue, seminal papers by Iyengar (2005) and Nilim & El Ghaoui (2005) independently show that a robust formulation with the $(s,a)$-rectangular assumption can be solved recursively via the robust Bellman equation (details provided later), thereby extending the results from non-robust DP theory. The rectangularity assumption allows the problem to be decomposed into independent subproblems. Therefore, they can adapt the value iteration and policy iteration for MDPs into robust counterparts, enabling efficient computation. This formulation is also referred to as robust DP and serves as a canonical form of RMDPs thereafter. Both papers introduce two imperative concepts. Firstly, they introduce uncertainty sets with statistical metrics, especially $\phi$-divergences, norms, and likelihood region ambiguity sets, which outperform polytopic models. Secondly, these works coincidentally emphasize that the rectangularity assumption is essential for the tractability of RMDPs, albeit without providing concrete proof. Bridging this theoretical gap, Bagnell et al. (2001) and Wiesemann et al. (2013) independently establish the NP-hardness of general (non-rectangular) RMDPs through distinct approaches.
While RMDPs have utilized statistical metrics to construct uncertainty sets and enhance representations of uncertainty, these methods do not incorporate prior distribution information about the uncertainty. Specifically, whether using an entropy metric or a likelihood region, the obtained uncertainty set only contains a collection of possible parameter realizations (i.e., each conditional transition probability) such that only their support is known. This is why such a collection is termed an uncertainty set. However, prior information is often available in practical applications, such as domain knowledge. Considering the potential conservatism of traditional robust approach (Thiele 2010, Delage & Mannor 2010) and advances in distributionally robust optimization (DRO) (Popescu 2007, Delage & Ye 2010, Goh & Sim 2010), Xu & Mannor (2012) first propose the framework of distributionally robust MDPs (DRMDPs), where the optimal decision is against the worst distribution among a collection termed as an ambiguity set. Compared to traditional RMDPs, the new framework more conveniently integrates statistical information to mitigate conservatism. Both Xu & Mannor (2012) and Yu and Xu (2015) demonstrate that DRMDPs can be reduced to expected standard RMDPs under certain assumptions, thus retaining tractability. These research findings have laid a solid theoretical foundation for RMDPs and enable the thriving development of the RMDP community.
[…]
@@
The remainder of the paper is organized as follows. In Section 2, we briefly introduce the preliminary of RMDPs. In particular, we summarize the definitions of rectangularity and provide a new proof for the NP-hardness of non-rectangular RMDPs. In Sections 3, 4, and 5, we introduce, respectively, three popular types of RMDPs: parametric RMDPs, moment-based RMDPs, and discrepancy-based RMDPs. In these sections, we focus on generic formulations and key theoretical advances. In Section 6, we investigate methodologies for modeling coupled uncertainty that violates rectangularity assumptions. Finally, we review some novel works in Section 7 that have sparked our interest beyond the traditional mini-max framework and conclude in Section 8.
@@ 2 Preliminaries
We denote by $[T]=\{1,2,...,T\}$ the set of positive running indices up to $T$, and use $\bigotimes$ to represent the Cartesian product. Let $\Xi:=\bigotimes_{t\in[T]}\Xi_{t}$ be the entire sample space, where $\Xi_{t}$ is the sample space at stage $t$. Let $\mathcal{F}$ be the $\sigma$-algebra of $\Xi$, and $\mathcal{F}$ consists of all subsets of $\Xi$. Exactly, the $\sigma$-algebra $\mathcal{F}$ is a set of events, where an event can be viewed as a scenario trajectory in the context of this paper. For the events that can be observed up to $t$, we denote the set of them by the filtration $\mathcal{F}_{t}$. Thus, we have $\left\{\emptyset,\Xi\right\}=\mathcal{F}_{0}\subseteq\mathcal{F}_{1}\subseteq\cdots\subseteq\mathcal{F}_{T}=\mathcal{F}$. One can think of this information structure also in terms of a scenario tree, $\mathcal{F}_{t}$ is generated by the partitions corresponding to nodes at stage $t$. Let $\mathfrak{M}(\Xi,\mathcal{F})$ be a set of probability measures on measurable space $(\Xi,\mathcal{F})$. We denote $\mathcal{P}\subseteq\mathfrak{M}(\Xi,\mathcal{F})$ by the set of probability measures we are interested in, and $\mathbb{P}\in\mathcal{P}$ is a probability measure on $(\Xi,\mathcal{F})$. Similarly, we specify $\mathcal{P}_{t}=\left\{\mathbb{P}(\cdot|\mathcal{F}_{t})\ |\ \mathbb{P}\in\mathcal{P}\right\}$ where $\mathbb{P}(\cdot|\mathcal{F}_{t})$ is the probability measure $\mathbb{P}$ conditioned on the filtration $\mathcal{F}_{t}$. In this survey, we particularly focus on the most common case that the (conditional) transition probabilities are time-invariant. This assumption is common in the literature, such as inventory and queueing systems, implying that system dynamics do not change over time. Except for otherwise noted, we assume stationary environments thereafter. As $\Xi$ is finite, without loss of generality, we can appropriately derive the corresponding probabilities $\mathbf{p}$, a (discrete) distribution $\mu$ or a probability measure $P$ on a single stage measurable space $(\Xi_{t},\mathcal{F}^{t})$ for all $t\in[T]$ with respect to $\mathbb{P}$, where $\mathcal{F}^{t}$ is the corresponding $\sigma$-algebra with respect to $\Xi_{t}$. Let $\Delta^{d}=\left\{\mathbf{p}\ |\ p_{1}+\cdots+p_{d}=1,\ p_{i}\geq 0\ \forall i\in[d]\right\}\in\mathbb{R}^{d}$ be the probability simplex, which represents the set of all valid probability distributions over $d$ outcomes, and $\bm{1}(\cdot)$ be an indicator function. We use superscript 00 to denote true underlying true value, e.g., $\mathbf{p}^{0}$, and $\hat{\cdot}$ denotes the estimated or nominal one, e.g., $\hat{\mathbf{p}}$.
As discussed in the introduction, we focus on sequential decision-making problems where the decision-maker observes the current state of the system, implements a feasible action, and stochastically transits to a new state with associated rewards. We consider a discounted case where the planning horizon $T$ can be infinite. Particularly, the formulations with $T=\infty$ often lead to the existence of stationary policies, which simplifies analysis and contributes to comprehending long-run behavior. When $T=\infty$, we usually require a discount factor $\gamma<1$ to ensure that the interested problems are well-defined. Let $\mathcal{S}$ denote the state space, and $\mathcal{A}$ denote the action space. We use a subscript $s$ to denote the value associated with the state $s$, e.g., $\mathcal{A}_{s}$. Here, we assume the state space and action space at each stage are time-invariant unless otherwise specified, i.e., $\Xi_{t}=\mathcal{S}\times\mathcal{A}$ for all $t\in[T]$. For $s\in\mathcal{S}$, $a\in\mathcal{A}$, let $p_{sas^{\prime}}:=P(s^{\prime}|s,a)$ denote the probability of transitioning to the next state $s^{\prime}\in\mathcal{S}$. These probabilities form the transition probability matrix $\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}$. When the transition $(s,a,s^{\prime})$ occurs, the decision-maker receives an immediate reward $R(s,a,s^{\prime})$. For simplicity, we assume $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is a known and deterministic mapping except for additional notes in the context. Given the discount factor $0<\gamma\leq 1$, such a multistage problem can be modeled as a MDP, which is represented as a six-element tuple $<T,\mathcal{S},\mathcal{A},R,\mathbf{p},\gamma>$. The goal of decision-makers is to maximize the cumulative discounted rewards by optimizing a policy $\pi$:
@@
where $R(s_{T+1})$ is the final reward function. A policy is a mapping from states to actions (deterministic) or a probability distribution over actions (stochastic). For example, given the state $s\in\mathcal{S}$, a deterministic policy $\pi(s)=a$ with probability 1, whereas a stochastic policy takes action $a$ with probability $\pi(a\mid s)$.
Once at a given state, we denote the value-to-go function $V_{t}(\cdot):\mathcal{S}\rightarrow\mathbb{R}$ as the mapping that quantifies the expected future reward generated by policy $\pi$ at stage $t$:
$V_{t}(s)=\max_{\pi}\mathbb{E}_{\mathbf{p}}^{\pi}\left[\sum_{\tau=t}^{T}\gamma^{\tau-1}R(s_{\tau},a_{\tau},s_{\tau+1})+\gamma^{T}R(s_{T+1})\,\middle|\,s_{t}=s\right].$ (2)
Fortunately, by leveraging the value-to-go function as given in (2), solving (1) is equivalent to recursively solving the following equations, known as the Bellman equations:
$V_{t}(s)=\max_{\pi}\mathbb{E}_{\mathbf{p}}^{\pi}\left[R(s_{t},a_{t},s_{t+1})+\gamma V_{t+1}(s_{t+1})\mid s_{t}=s\right],\ \forall\ t\in[T]$ (3)
where $V_{T+1}$ is the terminal value which takes the value $-R(s_{T+1})$ conventionally.
The Bellman equations decompose the multistage optimization problem into smaller subproblems by recursively relating the value of a state to the values of successor states. The advantages of solving the Bellman equation as in (3) are two-fold. First, the optimal equation retains the same form as the (optimal) Bellman equation, is that
$V^{*}(s)=\max_{\pi}\mathbb{E}_{\mathbf{p}}^{\pi}\left[R(s,\pi(s),s^{\prime})+\gamma V^{*}(s^{\prime})\right]$ (4)
Second, its recursive formulation brings significant computational benefits. For the finite horizon case, straightforwardly applying DP techniques, we can compute the optimal value via backward induction. For instance, in inventory management, backward induction can optimize ordering policies by considering expected immediate newsvendor costs and the value-to-go function at the next stage. For the infinite horizon case, we can compute the optimal value and policy via iterative methods, e.g., value iteration (VI) or policy iteration (PI) (Howard 1960). For example, we operate the maximization as on the right-hand side of (3) in VI and then update the value for each state. This process is repeated until the maximum deviation between two consecutive iterations is smaller than a predefined tolerance level $\epsilon$. Although the “curse of dimensionality” (Bellman 1966) is a known challenge, small to medium-scale MDPs can still be solved efficiently. For more detailed fundamentals, we refer to Puterman (2014).
Note that there are various representations of MDPs in different contexts, such as assuming known initial state distributions or continuous state spaces. The above definition of MDP captures the core components and remains widely applicable across this survey. Additional extensions will build on this formulation if necessary.
@@
The transition probability matrix $\mathbf{p}$ defined above is associated with the transition tuple $(s,a,s^{\prime})$, but plenty of literature assumes exogenous randomness. For example, in inventory management problems, the demand is often not affected by inventory position and order quantity. In this case, $\mathbf{p}$ may be independent of states and actions and it can be represented by a distribution directly.
@@
In this survey, we restrict our attention to MDPs with uncertain transition probabilities. To ensure generality, we consider the probability space $(\Xi,\mathcal{F})$, where each $\Xi_{t}=\mathcal{S}\times\mathcal{A}$, and the set $\mathcal{P}$ consists of probability measures of interest. However, to align with the literature and for simplicity, we also slightly abuse the notation $\mathcal{P}$ to represent the set of corresponding transition probabilities $\mathbf{p}$ or distribution $\mu$, which means $\mathcal{P}$ denotes the collection of uncertainty in this survey, including uncertainty set and ambiguity set. Formally, we consider RMDPs and DRMDPs, defined as follows:
@@
An RMDP is defined as a tuple $<T,\mathcal{S},\mathcal{A},\mathcal{P},R,\gamma>$, where the transition probability matrix $\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}$ is unknown but belongs to a known collection $\mathcal{P}$ of parameters, called uncertainty set. More precisely, RMDPs can be expressed as
@@
An DRMDP can be represented by a tuple $<T,\mathcal{S},\mathcal{A},\mathcal{P},R,\gamma>$, where the transition probability matrix $\mathbf{p}$ follows a distribution $\mu$, denoted by $\mathbf{p}\sim\mu$. The distribution $\mu$ is unknown but belongs to a known collection $\mathcal{P}$ of probability distributions, called the ambiguity set. More specifically, DRMDPs can be expressed as
@@
While both RMDPs and DRMDPs consider unknown transition probabilities, they handle this uncertainty in different ways. In RMDPs, the transition probability matrix $\mathbf{p}$ is regarded as an uncertain “parameter”, analogous to the uncertain outcomes in RO. The corresponding uncertainty set $\mathcal{P}$ contains all possible realizations of $\mathbf{p}$. The goal of (5) is to derive the optimal policy $\pi^{*}$ against the worst realization. Essentially, the only information available about the uncertainty in RMDPs is the support of $\mathbf{p}$, although this support can be refined using various methods, e.g., statistical metrics. In contrast, DRMDPs assume that additional prior distribution information about $\mathbf{p}$ is available. For instance, one might have a belief that an arbitrary transition probability $p_{sas^{\prime}}$ does not exceed 0.8 with at least 90% probability. Such belief can be readily incorporated into the ambiguity sets of DRMDPs, but not into the uncertainty sets of RMDPs. This distinction about how to treat transition probabilities is noteworthy, bringing various perspectives and enabling more powerful ambiguity set modeling. For example, moment-based ambiguity sets are more prevalent in DRMDPs compared to RMDPs.
Although RMDPs and DRMDPs differ in certain aspects, their formulations are closely aligned within the scope of this article, and the high-level modeling approaches exhibit substantial parallels. In fact, with appropriate modeling and practical considerations, one formulation might be transformed into the other (e.g., exogenous randomness specified later). Thus, this article mainly investigates the literature from a modeling framework perspective rather than emphasizing the distinctions between RMDPs and DRMDPs. Unless explicitly noted otherwise, RMDPs and DRMDPs will both be referred to as RMDPs in the following for simplicity.